OpenStudy (anonymous):

A treasure chest has a volume of x3 + 9x2 + 26x + 24 cm3 and the height is x + 4 cm. Find the polynomial that would represent the area of the bottom of the treasure chest? Explain your reasoning.

OpenStudy (experimentx):

x3+4x2+5x2+20x+6x+24 =x2(x+4)+5x(x+4)+6(x+4) =(x+4)(x2+5x+6) =(x+4)(x+3)(x+2) so answer would be (x+3)(x+2)

OpenStudy (anonymous):

ohmygod, thanks so much (: